OpenStudy (a1234):

Use the info given to evaluate the cosecant trig function, if possible. cos(pi/2 - x) = 3/5, cos x = 4/5 Is this 5/3? I don't get the question completely.

OpenStudy (a1234):

@peachpi

OpenStudy (anonymous):

\(cos (\pi/2 - x) = sin x\) → that's a co-function identity so they want the cosecant of the angle whose sine is 3/5 and cosine is 4/5

OpenStudy (anonymous):

so yeah, that's 5/3

OpenStudy (a1234):

Oh, ok. I forgot the co-function identity part. I just drew a triangle with the cos = 4/5 part.

OpenStudy (anonymous):

that's usually how I do it. and then just use the second given trig function to make sure the sign is correct

OpenStudy (a1234):

Yes. Thanks!